# Generalizable Deepfake Detection via Effective Local-Global Feature Extraction

### v1

@@ 4.3 Quantitative analysis
NPR 70.6 75.1 77.2 83.2 76.6 82.6 76.8 82.8 78.0 83.8 75.2 77.7 74.5 78.2 61.9 71.6 73.9 79.4
ours 80.2 83.7 80.8 93.0 73.0 79.4 78.4 84.1 77.8 87.9 83.2 88.7 75.5 78.0 71.3 89.6 77.6 85.6
Table 5: Cross-Diffusion-Sources Evaluation on the GenImage Dataset Zhu et al. (2024).
Method
Mean Acc. of
16 sub-testsets
CNNDetectionWang et al. (2020) 70.8
LGradTan et al. (2023) 75.3
OjhaOjha et al. (2023) 78.4
FreqNetTan et al. (2024a) 78.1
NPRTan et al. (2024b) 81.8
ours 84.7
Table 6: The mean accuracy of all 16 generation models on GenImage Zhu et al. (2024) dataset.
DWT Window Tiling FFT WAttention mean Acc. mean A.P.
$\surd$ 67.2 70.8
$\surd$ $\surd$ 49.5 52.3
$\surd$ $\surd$ 82.3 87.6
$\surd$ $\surd$ $\surd$ 83.9 90.2
$\surd$ $\surd$ $\surd$ $\surd$ 84.7 90.9
Table 7: Ablation study on the GenImage Zhu et al. (2024).
Conclusion. To comprehensively evaluate the superiority of our method, we test it on the entire GenImage dataset Zhu et al. (2024) and calculated the average accuracy. The GenImage dataset offers extensive coverage, including content generated by GANs as well as high-definition images from diffusion models. As shown in Table 6, our method achieves optimal detection performance, with an accuracy of 84.7%, significantly surpassing the accuracy of NPR and FreqNet, which are 81.8% and 78.1%, respectively.
@@
Figure 4: Logit Distributions of Extracted Forgery Features. We compare the state of the art NRP Tan et al. (2024b) and our method, both tuned with 4-class ProGAN Karras (2017) data. A total of four testing GANs and diffusion models are considered, including ProGAN Karras (2017), StyleGAN Karras (2019), IDDPM Nichol and Dhariwal (2021) and VQDM Gu et al. (2022), each randomly sampled 1k real and 1k fake images.
4.4 Qualitative Analysis
@@
4.5 Ablation Study
To thoroughly evaluate the effectiveness of each component in our method, we conduct ablation experiments by systematically removing some of the proposed modules. The results of these experiments, shown in Table 7, demonstrate the contribution of each module to deepfake detection. Notably, when we remove our main module (LoSFB), the performance of our method drops significantly, falling below 50%. This highlights the crucial role of local spatial-frequency domain feature extraction in detecting forged information.To further validate the effectiveness and rationale behind using the phase component of FFT to extract global information and employing sliding window tiling for local feature extraction in DWT, we remove these modules individually. As shown in the table, the accuracy decreased by 2.4% and 0.8%, respectively. These results emphasize the importance of both effectively extracting local texture information and subtle forged textures embedded in the global information.
5 Conclusion
In this paper, we propose a novel and effective method for deepfake detection, leveraging both local spatial-frequency information and global frequency domain features through two core modules: LoSFB and GloFB. This plug-and-play feature extractor is meticulously designed to capture a broader and more effective set of artifact features. Extensive experiments on 34 diverse generative models strongly demonstrate the generalization capability of our method.
6 Acknowledgments
This work is supported by the National Key R&D Program of China under grant 2021YFB2700900; by the National Natural Science Foundation of China under grant U22B2062, 62172232; by the Jiangsu Basic Research Programs-Natural Science Foundation under grant numbers BK20200039; by the Collaborative Innovation Center of Atmospheric Environment and Equipment Technology (CICAEET) fund.

### v2

@@ III-C Dual Frequency Branch Feature Enrichment Module
NPR 70.3 74.9 77.0 83.1 76.4 82.4 76.6 82.7 77.8 83.5 74.9 77.4 74.2 77.9 61.9 71.6 81.4 84.7
FatFormer 78.5 91.8 88.1 96.0 56.1 62.8 67.8 81.1 68.1 81.1 86.9 97.0 73.1 85.9 69.7 81.9 86.1 92.6
Ladeda 74.0 84.4 90.6 96.6 86.9 93.5 89.8 95.5 90.6 95.8 76.6 84.5 87.6 93.0 83.0 93.9 86.6 92.8
C2P_CLIP 77.2 95.8 88.5 98.8 59.2 83.0 82.8 97.3 82.8 97.2 87.2 98.4 80.1 95.5 65.1 92.9 88.2 92.6
AIDE 83.1 95.8 85.9 95.3 69.1 78.1 83.2 95.2 82.8 93.9 80.3 92.2 83.5 95.0 86.6 97.6 82.8 91.8
Ours 94.2 97.3 95.2 98.7 94.7 97.8 94.2 97.6 94.5 98.0 96.5 98.3 93.5 96.9 96.8 99.2 90.6 93.6
TABLE V: Cross-model Accuracy (Acc.) & Average Precision Score (A.P.) Performance on the AIGIBench Datasets [49].
Method ProGAN R3GAN StyleGAN3 StyleGAN-XL StyleSwim WFIR BlendFace E4S FaceSwap
CNNDetection 97.8 100.0 49.9 48.4 64.0 91.6 49.7 60.0 58.0 88.9 62.1 89.4 49.3 34.2 49.6 35.2 50.0 44.3
LGrad 98.0 99.9 65.6 69.1 59.9 64.3 61.2 62.8 68.1 75.0 55.6 53.1 38.6 39.4 39.2 39.4 28.2 33.3
CLIPDetection 99.5 100.0 93.0 98.6 68.4 79.0 87.1 95.0 94.5 99.3 91.0 97.8 47.2 42.7 72.4 82.5 69.3 79.4
FreqNet 99.3 100.0 62.3 56.8 83.0 92.4 79.8 84.1 80.8 91.8 58.5 48.9 23.3 34.1 25.8 34.7 40.4 43.4
NPR 99.8 100.0 80.8 75.6 84.1 89.4 74.5 75.9 84.7 91.7 77.8 77.0 33.9 33.3 34.0 34.4 41.1 39.5
FatFormer 99.9 100.0 94.4 99.3 91.5 96.9 93.4 98.0 93.9 99.8 88.2 98.5 46.0 44.7 70.0 73.0 77.6 83.4
LaDeDa 100.0 100.0 72.9 69.2 83.7 86.3 82.1 84.2 82.6 86.4 84.5 91.3 28.2 38.0 29.0 42.4 41.2 40.2
C2P_CLIP 100.0 100.0 95.4 99.8 93.8 98.5 94.1 98.1 95.2 100.0 94.8 99.5 55.2 60.5 69.8 76.1 73.5 82.2
AIDE 98.0 99.8 89.0 94.8 79.2 77.8 84.4 87.7 85.2 85.1 90.3 97.0 43.4 43.8 39.4 38.3 46.7 43.7
Ours 99.8 100.0 94.6 94.9 93.2 98.2 82.3 86.4 94.7 98.6 52.2 46.8 44.1 44.0 45.1 46.4 50.8 66.8
Method InSwap SimSwap FLUX1-dev Midjourney V6 GLIDE DALLE-3 Imagen3 SD3 SDXL
CNNDetection 50.1 44.3 49.7 45.3 50.3 50.3 49.9 39.3 51.5 67.3 50.4 53.8 51.0 50.3 52.8 66.3 52.6 65.5
LGrad 37.7 38.5 36.1 38.1 64.6 70.0 59.2 62.1 56.7 66.2 49.5 48.6 61.4 66.9 66.2 71.4 64.0 66.3
CLIPDetection 48.9 48.2 50.1 51.5 46.3 39.6 44.8 35.6 69.4 81.7 46.8 38.4 48.1 44.1 49.0 48.6 50.7 53.6
FreqNet 37.5 42.1 36.5 41.9 78.5 87.3 57.6 57.7 75.8 77.4 66.2 61.0 73.6 80.7 77.6 84.0 86.4 93.6
NPR 41.0 36.9 40.5 38.5 87.2 93.6 69.7 71.1 78.2 85.2 57.6 61.9 84.6 91.0 86.4 93.5 87.0 93.7
FatFormer 57.9 66.4 61.7 72.2 48.3 41.8 47.6 43.5 84.5 92.2 50.1 46.8 46.9 43.1 62.9 68.5 75.0 83.0
LaDeDa 38.3 41.8 37.3 39.6 84.2 88.0 67.0 68.0 86.4 89.4 39.2 46.3 81.1 84.6 83.6 87.6 83.1 88.0
C2P_CLIP 59.5 69.0 67.9 78.5 47.6 46.8 51.2 50.2 85.1 93.4 64.4 75.4 45.2 46.5 53.5 65.6 59.3 76.0
AIDE 46.1 43.0 47.2 46.8 84.9 86.6 77.9 78.1 90.0 89.3 56.9 57.0 86.4 90.8 89.7 94.1 88.9 90.0
Ours 52.1 63.7 48.4 57.7 93.2 96.8 86.4 92.7 94.2 98.4 46.2 49.0 93.6 96.0 95.4 97.6 95.9 98.7
Method BLIP Infinite-ID InstantID IP-Adapter PhotoMaker SocialRF CommunityAI Mean
Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P. Acc. A.P.
CNNDetection 49.8 44.7 50.8 55.8 50.1 67.1 50.0 50.1 49.6 51.5 50.3 46.3 49.9 50.5 53.6 57.6
LGrad 56.4 57.8 31.3 34.4 53.9 54.8 45.5 46.3 32.0 35.8 49.6 49.3 56.4 63.0 53.4 56.2
CLIPDetection 57.8 68.9 72.4 84.2 67.4 79.5 62.3 73.9 47.0 41.4 49.2 46.2 49.9 50.0 63.3 66.4
FreqNet 93.8 99.9 79.0 74.5 79.8 86.3 78.8 79.9 77.0 74.9 54.2 58.1 55.9 69.7 60.2 61.7
NPR 94.0 97.0 62.0 62.8 71.3 75.9 79.0 81.1 46.4 52.0 54.5 58.4 54.4 60.0 68.2 70.8
FatFormer 82.8 92.3 83.3 91.2 69.3 77.4 67.1 74.1 46.4 37.4 56.9 65.1 51.9 69.9 69.9 74.3
LaDeDa 93.4 96.7 79.0 67.5 81.0 81.8 82.3 82.2 69.3 66.6 56.2 61.7 54.8 55.9 68.8 71.3
C2P_CLIP 84.9 93.8 85.2 93.4 86.8 93.9 55.1 65.9 50.6 52.0 56.4 62.3 51.0 54.5 71.0 77.3
AIDE 92.7 96.7 88.9 87.9 87.6 89.0 87.1 91.0 83.4 80.9 56.8 62.1 51.3 52.8 74.9 76.2
Ours 99.2 99.7 90.0 91.0 94.6 97.7 93.3 96.6 92.7 95.2 60.8 70.8 54.6 66.6 77.9 82.0
IV Experiments
To verify the excellent AI-generated image detection performance of our method, we follow the paradigms of [43] and [50], which will be described in detail below.
IV-A Dataset
Training set. To ensure a consistent basis for comparison, we use the training set of ForenSynths[50]. The training set consists of 20 different categories, each containing 18,000 synthetic images generated using ProGAN, and an equal number of real images from the LSUN dataset. Following the paradigms of [43], we adopt specific 4-class training settings, denoted as (car, cat, chair, horse).
Testing set. In order to comprehensively compare the effectiveness of our proposed method, we select different GAN and diffusion model-based generated datasets for testing, which contain a wide range of real-world images.
  • GANGen-Detection [9]: To evaluate more realistic scenarios, we extend our evaluation using 9 additional GAN-generated images. There are 4K test images for each model, with equal numbers of real and fake images.
  • DiffusionForensics [46]: To expand the testing scope, we adopt the diffusions dataset of DIRE [46] for evaluation, including ADM [23], DDPM [51], IDDPM [52], LDM [24], PNDM [53], VQ-Diffusion [54], Stable Diffusion v1 [24], Stable Diffusion v2 [24]. The real images are sampled from LSUN and ImageNet datasets.
  • UniversalFakeDetect [43]: This test set contains images generated from Glide, DALLE-mini [55], LDM [24]. It adopts images of LAION and ImageNet datasets as the real data.
  • AIGCDetectBenchmark [48]: This dataset comprehensively includes GAN and diffusion, and introduces some new diffusion generation methods as detection datasets, including ProGAN, StyleGAN, StyleGAN2, BigGAN, CycleGAN, StarGAN, GauGAN, Deepfake, WFIR, Midjourney, SDv1.4, SDv1.5, ADM, GLIDE, Wukong, VQDM, which is a huge challenge for deepfake detection.
  • AIGIBench [49]: AIGIBench is the first benchmark to conduct a comprehensive evaluation of four critical components in the AIGI detection pipeline, and comprehensively simulates state-of-the-art image generation methods, including: (a) GAN-based noise-to-image generation (ProGAN, StyleGAN3, StyleGAN-XL, StyleSwim, R3GAN, and WFIR), (b) Diffusion for text-to-image generation (SD-XL, SD-3, DALLE-3, Midjourney-v6, FLUX.1-dev, Imagen-3, and GLIDE), (c) GANs for deepfake (BlendFace, E4S, FaceSwap, InSwap, and SimSwap), and (d) Diffusion for personalized generation (InstantID [27], Infinite-ID [28], PhotoMaker [29], BLIP-Diffusion [30], and IP-Adapter [31]). It also includes 2 general subsets featuring fake images collected from social media platforms, which called CommunityAI and SocialRF.
IV-B Implementation Details
We design a lightweight CNN network using convolutional layers and Resnet network as classifiers. The network is trained using the Adam optimizer[56] with a learning rate of $2\times 10^{-4}$. The batchsize is set to 128, and we train the network for 90 epochs. We also use a learning rate decay strategy that reduces the learning rate by twenty percent after every ten epochs. What’s more, the hyperparameter $\lambda$ is set to 0.4. For the loss function, we used BCELoss, which is commonly used in classification tasks. Our approach is implemented using PyTorch on an NVIDIA RTX A6000 GPU. To evaluate the performance of the proposed method, we follow the evaluation metrics used in baselines [43, 9], which include average precision score (A.P.) and accuracy (Acc.).
IV-C Quantitative analysis
We compare with the previous methods: CNNDetection [50], Patchfor [34], F3Net [37], FrePGAN [7], GANDetection [57], LGrad [8], CLIPDetection [43], FreqNet [9] ,NPR [3], FatFormer [15], Ladeda [35], C2P_CLIP [17], AIDE [18]. We conduct comprehensive experiments on five datasets to verify the effectiveness of our method. Specifically, the results of Patchfor [34], F3Net [37], FrePGAN [7], GANDetection [57] all come from NPR [3], and we obtain results of FatFormer [15], C2P_CLIP [17] using the official pre-trained model. In addition, we retrain all other methods following the paradigms of [43], which adopts specific 4-class training settings, denoted as (car, cat, chair, horse).
GANGen-Detection[9]. Since our method uses ProGAN for training, we first perform cross-domain testing on GAN. As shown in Table I, our detection on GAN has achieved good results. Specifically, it achieves mean Accuracy (Acc.) and mean Average Precision (A.P.) of 97.3% and 98.8%, respectively. Compared to Ladeda and C2P_CLIP, our method improves mean Acc. by 0.1% and 0.9%, respectively. However, on Fatformer, the accuracy is 1.1% higher than our method, which needs to be improved.
DiffusionForensics [46] & UniversalFakeDetect [43]. To further evaluate the generalization ability of our method, we train on ProGAN and test on images generated by various diffusion models. The results are presented in Tables II and III. Despite being trained exclusively on ProGAN-generated images, our method demonstrates strong generalization capabilities across diverse diffusion models. Specifically, it achieves mean Accuracy (Acc.) and mean Average Precision (A.P.) of 98.8% and 100.0%, respectively. Compared with the best detection methods NPR and Ladeda, our method improves mean Acc. by 3.0% and 3.7%, respectively. However, methods Fatformer, C2P_CLIP, and AIDE, which have good results on other datasets, have an accuracy rate of less than 90% on this dataset, and even C2P_CLIP has only 80.7% accuracy. The most direct speculation for this result is that they used CLIP in the network skeleton, but the real reason is still worth further exploration. Additionally, on the UniversalFakeDetect dataset [43], our method achieves an average accuracy of 96.4%, 2.2% accuracy improvement over the best method.
GenImage [48]. Compared with testing only on diffusion or GAN, GenImage not only includes the classic GAN generation method, but also adds new and more realistic diffusion generation datasets, such as DALLE2. The test results are shown in Table IV. Our method achieves a notable improvement in mean accuracy, outperforming C2P_CLIP and Ladeda by 2.4% and 4.0%, respectively, and reaching an accuracy of 90.6%.
AIGIBench [49]. AIGIBench proposes a more challenging forgery detection dataset to evaluate the performance of forgery detection models from different perspectives. The test results are shown in Table V, our detection has achieved good results. Specifically, it achieves mean Accuracy (Acc.) and mean Average Precision (A.P.) of 77.9% and 82.0%, respectively. Compared to Ladeda and C2P_CLIP, our method improves mean Acc. by 3.0% and 5.9%, respectively. Although our results are the best, there are still major problems and improvements: 1) Although our method has achieved an accuracy rate of more than 90% in most GAN and diffusion scenarios, it still has an accuracy rate of only about 60% in the dataset collected in unknown generated social scenarios. Other methods also have very low accuracy rates, which still needs further research and improvement. 2) When it comes to face-swapping forgery detection, the performance is generally low, even less than 50%. Almost all methods have this problem, which is also a problem that needs to be considered and improved.
TABLE VI: Module Ablation Studies on the Five Datasets.
DWT Window Tiling FFT RSWAttention GANGen-Detection DiffusionForensics UniversalFakeDetect GenImage AIGIBench mean Acc. / A.P.
49.6 / 49.6 48.6 / 48.6 47.9 / 47.9 49.4 / 49.4 49.6 / 49.6 49.0 / 49.0
93.2 / 96.5 95.6 / 100.0 93.2 / 95.3 87.5 / 91.2 74.1 / 78.5 88.7 / 91.2
95.2 / 97.1 96.9 / 100.0 95.0 / 97.4 88.7 / 92.5 75.5 / 79.8 90.3 / 93.4
96.2 / 98.8 96.6 / 99.9 96.2 / 99.1 90.0 / 92.1 74.2 / 75.8 90.6 / 93.1
96.6 / 98.8 98.8 / 100.0 96.4 / 99.3 90.6 / 93.6 77.9 / 82.0 92.1 / 94.1
TABLE VII: RSWAttention Validity Studies on the Five Datasets.
Attention GANGen-Detection DiffusionForensics UniversalFakeDetect GenImage AIGIBench mean Acc. / A.P.
WAttention 70.7 / 82.9 65.8 / 89.5 73.2 / 88.4 71.5 / 81.2 58.3 / 63.2 67.9 / 81.0
RSWAttention 96.6 / 98.8 98.8 / 100.0 96.4 / 99.3 90.6 / 93.6 77.9 / 82.0 92.1 / 94.1
TABLE VIII: Phase-only Part of FFT Validity Studies on the Five Datasets.
FFT GANGen-Detection DiffusionForensics UniversalFakeDetect GenImage AIGIBench mean Acc. / A.P.
Phase Amp
95.2 / 97.1 96.9 / 100.0 95.0 / 97.4 88.7 / 92.5 75.5 / 79.8 90.3 / 93.4
96.6 / 98.8 98.8 / 100.0 96.4 / 99.3 90.6 / 93.6 77.9 / 82.0 92.1 / 94.1
93.0 / 97.0 91.3 / 99.2 92.2 / 97.2 85.1 / 91.2 72.7 / 77.8 86.9 / 92.5
TABLE IX: DWT Different Frequency Domain Subbands Studies on the Five Datasets.
DWT GANGen-Detection DiffusionForensics UniversalFakeDetect GenImage AIGIBench mean Acc. / A.P.
LL 65.7 / 79.9 63.6 / 84.7 68.5 / 82.2 66.0 / 78.0 57.3 / 62.3 64.2 / 77.4
LH 75.9 / 89.3 68.3 / 87.0 74.3 / 89.1 70.4 / 82.3 60.2 / 66.0 69.8 / 82.7
HL 78.9 / 88.3 69.7 / 86.2 74.5 / 85.6 71.8 / 82.7 61.7 / 66.2 71.3 / 81.8
HH 77.3 / 88.7 67.2 / 87.2 74.6 / 85.9 71.9 / 84.6 61.5 / 67.7 70.5 / 82.8
HL,HH 91.8 / 96.3 81.4 / 99.3 91.6 / 97.7 81.4 / 90.3 71.5 / 76.7 80.3 / 88.0
LH,HL,HH 90.7 / 98.0 84.0 / 99.9 92.5 / 99.0 84.3 / 93.0 73.8 / 80.5 85.1 / 94.0
LL,LH,HL,HH 96.6 / 98.8 98.8 / 100.0 96.4 / 99.3 90.6 / 93.6 77.9 / 82.0 92.1 / 94.1
TABLE X: Hyperparameter $\lambda$ Ablation Experiment on the Five Datasets.
$\lambda$ GANGen-Detection DiffusionForensics UniversalFakeDetect GenImage AIGIBench mean Acc. mean A.P.
0.0 95.2 / 97.1 96.9 / 100.0 95.0 / 97.4 88.7 / 92.5 75.5 / 79.8 90.3 93.4
0.1 95.7 / 98.9 97.2 / 100.0 95.8 / 99.0 89.4 / 93.0 77.0 / 81.1 91.0 94.4
0.2 96.2 / 98.8 97.5 / 100.0 95.3 / 98.9 90.0 / 93.2 77.0 / 81.6 91.2 94.5
0.4 96.6 / 98.8 98.8 / 100.0 96.4 / 99.3 90.6 / 93.6 77.9 / 82.0 92.1 94.7
0.6 95.5 / 98.9 98.5 / 99.9 96.0 / 99.0 89.9 / 93.1 76.9 / 80.9 91.1 94.4
0.8 95.0 / 97.8 93.4 / 100.0 93.7 / 98.5 89.9 / 94.2 76.5 / 80.2 89.7 94.1
1.0 49.6 / 49.6 48.6 / 48.6 47.9 / 47.9 49.4 / 49.4 49.6 / 49.6 49.0 49.0
@@ IV-C Quantitative analysis
Figure 4: Logit Distributions of Extracted Forgery Features. We compare the state of the art NRP [3] and our method, both tuned with 4-class ProGAN [19] data. A total of four testing GANs and diffusion models are considered, including ProGAN [19], R3GAN [58], BLIP [30] and VQDM [54], each randomly sampled 1k real and 1k fake images.
IV-D Qualitative Analysis
@@
IV-E Module Ablation Studies.
In order to comprehensively evaluate the effectiveness of our method, we conduct ablation studies on each module. The results of these experiments, shown in Table VI, demonstrate the contribution of each module to AI-generated image detection. Specifically, the experiment can prove that:
FFT alone has negative impacts on Attention. When we remove our main module (DWT-Based Window Tiling), the performance of our method drops significantly, falling below 50%. In our design, FFT is integrated as a parallel branch to complement DWT, leveraging its phase information to capture long-range dependencies that enhance fine-grained feature richness for reconstructed sliding window attention. However, the localized focus of RSWAttention, optimized for fine-grained details, appears incompatible with FFT’s global frequency representation when not moderated by DWT’s multi-scale decomposition or Window Tiling’s spatial structuring. This misalignment suggests that FFT alone introduces overly coarse features, diluting the attention mechanism’s ability to discern subtle, localized artifacts critical for deepfake identification.
Reasonable design of sliding window greatly improves the ability of attention extraction artifacts. Whether it is DWT or window tiling, the lack of any part will lead to a decrease in model accuracy. Without DWT, window attention cannot capture fine-grained artifacts. Removing Window Tiling yields a mean Acc. / A.P. of 90.3% / 93.4%, a notable decline from the full model’s 92.1% / 94.1%, suggesting that Window Tiling preserves critical spatial relationships within local windows, allowing RSWAttention to exploit Dependencies between local features. Without it, the disruption of these relationships impairs the model’s ability to detect localized artifacts, which is very important in datasets with complex or subtle manipulations. Therefore, carefully designed in-window features based on DWT and window tiling are crucial for RSWAttention to extract fine-grained forgery features.
It is effective to use the phase part of FFT as another branch.As shown in Table VI, when the phase part of FFT is excluded, the average accuracy of the model decreases 1.8%, indicating that FFT enhances the attention mechanism’s capacity to model interdependencies between distant image regions. FFT, as another frequency domain branch, enriches the local features within the window attention from different perspectives.
IV-F Comparative Studies of Module Design
To further verify that our proposed method helps the model extract more fine-grained multi-angle comprehensive forgery traces while modeling the importance and implicit dependencies between local features, we conducted further comparative studies on the modules, as follows.
1) The reconstructed sliding window attention mechanism greatly improves the ability of attention to capture forgery features. As shown in Table VII, compared with WAttention, our RSWAttention improves the mean Acc. and mean A.P. by 24.2% and 13.1% respectively, which suggests that the RSWAttention greatly promotes the attention mechanism to capture fine-grained forgery traces within the local window while modeling the importance and interdependence between local features.
2) Utilizing only the phase component of the FFT can reduce features not related to forgery compared to using the full FFT. As shown in Table VIII, if we use not only the phase part of the FFT but also the amplitude part as input to the network, the average accuracy of the model drops by 5.2%, confirming that amplitude introduces noise or irrelevant intensity information that confounds the detector. In contrast, the phase, encoding structural details like edges and shapes, aligns with the artifact-relevant features needed for detection. This finding corroborates prior image processing research, where phase information is prioritized over amplitude for structural analysis, reinforcing our design choice to leverage FFT’s phase as a complementary input branch.
3) All subbands of DWT are indispensable. Each frequency domain subband of DWT has its own unique meaning, the low-frequency part retains the overall structure and semantic background of the image, and the high-frequency part captures subtle texture and edge information. Deleting any subband will reduce the comprehensiveness of the information in the window. As shown in Table IX, whether using only one frequency domain subband or omitting some frequency domain subbands, the performance of the model will be degraded.
IV-G Hyperparameter $\lambda$ Ablation Study.
In order to more comprehensively evaluate the effectiveness of our method, we also conduct ablation experiments on the hyperparameter $\lambda$, which can be shown in Table X. When our parameter $\lambda$ is set to 0.4 as set in the paper, the best results are achieved on all datasets.
V Conclusion
In this paper, we propose a novel and effective method for AI-generated image detection. Specifically, we designed a reconstructed sliding window attention mechanism, which reconstructs the features within the window and limits the attention to the local window, forcing the attention to extract fine-grained features while modeling the importance and dependencies between internal elements in the local area. In addition, we design a dual frequency branch framework consisting of four frequency bands in the DWT domain and the phase part in the FFT domain to extract richer local artifact features from multiple perspectives. Extensive experiments on 65 diverse generative models strongly demonstrate the generalization capability of our method.